OpenStudy (anonymous):

whats the maclaurin series of fx = (1-cos(x^7)/ x^3)....? need some help ...

OpenStudy (anonymous):

thats pretty easy

OpenStudy (anonymous):

the denominator is already in powers of x, leave that alone

OpenStudy (anonymous):

then just need to know the general maclaurin series for cos(x)

OpenStudy (anonymous):

\[\cos(x) = 1 - \frac{x^2}{2!} + \frac{x^4}{4!} - .... \]

OpenStudy (anonymous):

replace the x with x^7

OpenStudy (anonymous):

\[(1- \cos(x^7) ) = (1 - ( 1- \frac{(x^7)^2}{2!} + \frac{(x^7)^4}{4!} - ..... ) )\]

OpenStudy (anonymous):

\[(1-\cos(x^7) ) = \frac{x^{14}}{2!} - \frac{x^{28}}{4!} + .... \]

OpenStudy (anonymous):

\[\frac{1-\cos(x^7)}{x^3} = \frac{x^{11}}{2!} - \frac{x^{25}}{4!} + .... \]

OpenStudy (anonymous):

etc etc, you can get a general form if you really want,

OpenStudy (anonymous):

how do you know when to stop the serioes? the general form?

OpenStudy (anonymous):

it doesnt stop obviously

OpenStudy (anonymous):

you just write down the first two or three terms, then write .... after wards

OpenStudy (anonymous):

so they know it continues , its an infinite series

OpenStudy (anonymous):

so if n=1 you just give em the first term? because they want the interval of convergence also